### v1

# Supersonic and Superluminal Energy and Speed of Information via Temporal Interference in a Dispersionless Environment

Abstract
Numerical implementation of a theory yields acoustic wave packets whose peak-to-peak speeds, $c_{3d}$, are supersonic in a dispersionless medium due to temporal interference between direct and boundary-reflected paths. The effect occurs when the source and receiver are near each other and at least one is within $c\tilde{\delta t}/2$ of the boundary, where $c$ is the phase speed of propagation in the medium, and $\tilde{\delta t}$ is the smallest temporal separation between the paths at which interference first occurs. This direct+reflected path effect is distinct from previously-observed superluminal occurrence of group speeds due to anomalous dispersion, quantum tunneling, and cavity vacuum fluctuations. For temporally interfering direct+reflected paths, simulations yield a speed of information less than $c$, though no proof exists $c$ cannot be exceeded. The speed of information from the interfering paths can exceed the speed derived from propagation only along the direct path. We conjecture these results will also hold for electromagnetic (EM) wave propagation, and so the effect may not violate special relativity. These theoretical and simulation results, as well as their conjectured EM extension, should be readily accessible to experimental verification.
I Introduction
Recent theoretical results for dispersionless media predict acoustic wave packets can occur at locations with speeds significantly slower or faster than the phase speed when the source and receiver are near each other, are near a reflecting boundary, and the waveforms from the direct and reflected paths temporally interfere [78, 85] (Fig. 1). In other words, a wave packets’ speed is modified even when the phase and group speeds are identical. For the direct+reflected path effect, the so-called $c_{3d}$ speed of a wave packet is defined as,
$c_{3d}\equiv l_{1}/t_{m}\ ,$ (1)
where $l_{1}$ is the distance of the straight path from source to receiver, and $t_{m}$ is the measured time between the peaks of the wave packets. The $c_{3d}$ can be unequal to the group speed. Unless noted otherwise, in this paper the wave packet peak is defined to coincide with the first arriving peak of the absolute value of the Hilbert transform of the time series. This slowing and speeding of wave packets may need to be accounted for when deriving correct confidence intervals of location based on measurements of propagation time or measurements of the time-differences of arrival (TDOA) between receivers [78, 85]. The physics of the direct+reflected path effect only requires a reflecting boundary with a nearby source or receiver, suggesting the same phenomenon might occur for electromagnetic (EM) waves near a mirror. Because the effect yields supersonic values of the $c_{3d}$ in dispersionless media, we conjecture it may yield superluminal values of the $c_{3d}$ for EM waves propagating in a vacuum near a reflecting boundary. This possibility is discussed in the context of special relativity [24] wherein the speed of information may not exceed the speed of light in a vacuum [60, 6, 7].
Figure 1: Signal propagates from source $\bf s$} to receiver $\bf r$} along direct and boundary reflected paths with lengths, $l_{1}$ and $l_{2}$, and propagation times, $t_{1}$ and $t_{2}$, respectively. The $y$ axis is not shown but the xyx-y plane is perpendicular to the $z$ axis and the boundary is in a xyx-y plane.
It has long been known superluminal group speeds occur due to anomalous dispersion [34], where the frequencies of an EM wave packet interact with charged particles near their resonant frequencies. In order to preserve the idea of causality in special relativity [24], the second postulate was modified to mean the speed of information cannot exceed the speed of light in a vacuum [60, 6, 7, 16]. Scientists have observed superluminal phenomena due to several physical effects including anomalous dispersion [16, 88] microwave tunneling [25, 45], quantum tunneling [9, 83], and discussed the possibility in the context of quantum electrodynamics [58]. None of these investigations yielded an observed speed of information exceeding the speed of light in a vacuum.
Consequently, the speed of information due to the direct+reflected path effect is simulated in this paper for acoustic waves following methods from information theory employed by Stenner et al. [84]. Whether Stenner et al. [84] actually measured superluminal energy was discussed shortly thereafter [50]. Regardless of this interpretation, the implications of our simulations for the speed of information are then extrapolated for application to EM waves.
Robertson et al. [55] indirectly inferred the existence of superluminal propagation of acoustic pulses in a short loop tube attached to a long tube. The loop tube introduced other acoustic paths, destructively and constructively interfering with the otherwise undisturbed signal at certain resonant frequencies. This mimicked anomalous dispersion, and caused the group speed to increase. An acoustic speaker broadcasted narrowband acoustic waves, near a resonant frequency of the loop, at one end of a straight tube of length eight meters. The propagation time was derived from data collected at a microphone at the other end of the eight meter tube. Times of the arriving acoustic energy were made with and without the loop. The arriving energy packet was advanced by 0.0024 s when the loop was present. Assuming a sound speed in air of 330 m/s, the speed of the energy packet in the absence of resonances is about $8\ \mbox{m}/330\ \mbox{(m/s)}\sim 0.024$ s, Thus, the propagation time of the wave packet with the loop is 0.024 - 0.0024=0.0218 s, yielding a measured group speed of $8\ \mbox{m}/0.0218\ \mbox{s}\sim 367$ m/s, much less than the speed of light in a vacuum. However, they explain there must be superluminal speeds in the proximity of the loop to cause this advance. They did not attempt to directly measure superluminal speeds because they believed placing the speaker and microphone directly at the start and end of the loop would destroy the natural resonances of acoustic waves in the loop and suppress the superluminal effect. Interestingly, Robertson et al. [55] state their phenomenon is like the so-called “Comb” effect from the field of architectural acoustics, wherein sound arriving at a listener is a combination of the direct path and one reflecting from a hard boundary, causing destructive and constructive interference at the listener’s head at certain frequencies, and degrading the music’s fidelity [56]. The direct+reflected path effect seems to be the next step, recognizing interference of direct and reflected paths modifies the $c_{3d}$, i.e. the speed of appearance of acoustic wave packets.
II Exact solutions for acoustic waves
Spiesberger et al. [78] recalled the exact three-dimensional solutions to the linear acoustic wave equation in the presence of an ideal flat boundary,
$\rho_{a}(x,y,z,t)$ $=$ $a_{1}s(t-l_{1})-a_{2}s(t-l_{2})\ ,$ (2)
$\rho_{b}(x,y,z,t)$ $=$ $a_{1}s(t-l_{1})+a_{2}s(t-l_{2}))\ ,\$(3)
where $\rho_{a}(x,y,z,t)$ is for a boundary with zero fluctuations of pressure and $\rho_{b}(x,y,z,t)$ is for zero normal velocity. When energy spreads spherically from the sources, $a_{i}=1/l_{i}$, where the distances, $l_{1}$ and $l_{2}$ are $l_{1}=\sqrt{(x_{s}-x_{r})^{2}+(y_{s}-y_{r})^{2}+(z_{s}-z_{r})^{2}}$ and $l_{2}=\sqrt{(x_{s}-x_{r})^{2}+(y_{s}-y_{r})^{2}+(z_{s}+z_{r})^{2}}$ with source and receiver at $(x_{s},y_{s},z_{s})$ and $(x_{r},y_{r},z_{r})$ respectively in Cartesian coordinates (Fig. 1). These solutions are most easily understood by introducing an image source on the opposite side of the boundary. It is $180^{\circ}$ out of phase for the boundary with zero pressure fluctuation and in phase for the boundary with zero normal velocity. $s(t)$ is the emitted waveform.
III Speed of Information
Stenner et al. [84] state they measured the speed of information for EM waves where the group speed exceeded the speed of light in a vacuum. They transmitted two symbols and found through experiment the speed of information was less than the speed of light in a vacuum. Their procedures are duplicated here to theoretically derive the speed of information when the $c_{3d}$ exceeds both the phase and group speeds, $c$, due to temporal interference between the direct and reflected paths.
Symbols 0 and 1 are detected at the receiver, where they are identical up to time $t=0$, defined to be the time when a switch is thrown to choose symbol 0 or 1 thereafter (Fig. 2). The point of non-analyticity occurs at $t=0$, and in our formulation there is a discontinuity in the first derivative of the ideal transmitted waveform. Both symbols have carrier frequency, $f_{s}$, but are distinguished by different envelopes. Before the switch is thrown, the waveform is,
$w_{a}(t)=\cos(2\pi f_{s}t)\exp[-(t/\tau_{a})^{2}]\ ;\ t\leq 0\ .$ (4)
The envelope for symbol 0 goes down for $t>0$ as,
$w_{0}(t)=\cos(2\pi f_{s}t)\exp[-(t/\tau_{b})^{2}]\ ;\ t>0\ .$ (5)
The envelope for symbol 1 goes up at $t=0$ to maximum amplitude, $a_{max}$, at $t=t_{max}$ and then down afterwards as,
$w_{1}(t)$ $=$ $a_{max}\cos(2\pi f_{s}t)\exp[-((t-t_{max})/\tau_{c})^{2}]\ ;\ 0<t\leq t_{max}$ (6)
$w_{1}(t)$ $=$ $a_{max}\cos(2\pi f_{s}t)\exp[-((t-t_{max})/\tau_{d})^{2}]\ ;\ t_{max}<t\ .$ (7)
where
$t_{\max}=\tau_{b}\sqrt{\ln(a_{max})}\ .$ (8)
See Fig. 2 for an example.
Figure 2: Envelopes of ideal symbols following Eqs. 4 through 8. Point of non-analyticity is time zero. Symbols 0 and 1 differ only at times exceeding zero with symbol 0 being dashed.
III.1 Symbol classification
The speed of information is measured from the time the switch is thrown to the time the symbol is reliably detected. Let $R_{j}(t)$ be the time series at the receiver for symbols $j=0\ \mbox{and}\ 1$ in the absence of noise. The symbols can be classified without error at the least time they differ, i.e. $T_{differ}$. This time decreases with bandwidth, and vanishes for infinite bandwidth filters.
To mimic reality, the switch and the receiver’s detector have finite bandwidth, each with their own time delay. The bandwidth’s are made large to minimize their delays. Following Stenner et al. [84] and principles of signal detection (pp. 88-94 of Madhow [37]), symbol classification at finite SNR becomes a statistical problem employing the decision statistic,
$D(\tau,\chi)\equiv{L}_{0}(\tau,\chi)-{L}_{1}(\tau,\chi)\ ,$ (9)
where $L_{j}(\tau,\chi)$ are outputs of a matched filter operating on the received time series,
$\chi_{j}(t)=R_{j}(t)+n(t)\ ,$ (10)
where $n(t)$ is noise, and the matched filters for symbols 0 and 1 are,
$L_{0}(\tau,\chi)$ $=$ $\int_{t_{s}}^{t_{s}+\tau}\chi_{j}(t)R_{0}(t)dt/[\alpha_{0}(\tau_{\alpha})N_{0}(\tau)]\ ,$ (11)
$L_{1}(\tau,\chi)$ $=$ $\int_{t_{s}}^{t_{s}+\tau}\chi_{j}(t)R_{1}(t)dt/[\alpha_{1}(\tau_{\alpha})N_{1}(\tau)]\ .$ (12)
The denominators are normalizing terms defined as,
$\alpha_{0}(\tau_{\alpha})$ $\equiv$ $\int_{t_{s}}^{t_{s}+\tau_{\alpha}}\chi^{2}_{0}(t)/N_{0}(\tau_{\alpha})\ ,$ (13)
$N_{0}(\tau)$ $\equiv$ $\int_{t_{s}}^{t_{s}+\tau}R_{0}^{2}dt\ ,$ (14)
$\alpha_{1}(\tau_{\alpha})$ $\equiv$ $\int_{t_{s}}^{t_{s}+\tau_{\alpha}}\chi^{2}_{1}(t)/N_{1}(\tau_{\alpha})\ ,$ (15)
$N_{1}(\tau)$ $\equiv$ $\int_{t_{s}}^{t_{s}+\tau}R_{1}^{2}dt\ .$ (16)
The matched-filter is integrated starting at time $t_{s}$, chosen to occur before the point of non-analyticity occurs at the receiver, so as to accumulate its energy between it and the ending evaluation time at $t_{s}+\tau$, where $\tau\geq t_{s}$ and continues on until the energy of the time series has passed. $\tau_{\alpha}$ is chosen to be just before the time when the symbols separate. For experiments, the replicas, $R_{j}(t)$, are estimated by averaging over many realizations to suppress noise. In this paper, they are equal to the received waveforms without noise.
To estimate the bit error rate (BER), the probability density function of $D(\tau,\chi)$ is empirically derived for two scenarios. In the first, it is empirically estimated from many noisy realizations of only symbol 0; i.e. many realizations of $\chi_{0}(t)$. The second probability distribution of $D(\tau,\chi)$ is derived from many noisy realizations of only symbol 1. Each probability density function is fitted to a Gaussian distribution, and normalized to area of one-half. Their area of overlap is the BER [84].
We verified our implementation of the BER empirically as follows. First, the noiseless values for $D(\tau,\chi)$, denoted $\hat{D}_{0}(\tau,\chi_{0})$ and $\hat{D}_{1}(\tau,\chi_{1})$ for symbols 0 and 1, are constructed separately for each symbol using the noiseless timeseries at the receiver. The hat indicates evaluation with no noise, i.e. $n(t)=0$. Then, for any incoming noisy symbol at time $\tau$, $D(\tau,\chi)$ is computed. Symbol 0 is declared to be received when $D(\tau)$ is closer to $\hat{D}_{0}(\tau,\chi_{0})$. Otherwise symbol 1 is declared received. The BER rate from this procedure agrees with the procedure outlined by Stenner et al. [84].
III.2 BER example
Consider an ideal scenario in water where the phase speed, $c$, is 1500 m/s and a compact source and receiver are located at Cartesian $(x,y,z)$ coordinates (0, 0, -11.746) m and (12.698, 0, -41.995) m, respectively. The carrier and sample frequencies are 3948.34 Hz and 100 kHz respectively. For the ideal signal, $\tau_{a}$, $\tau_{b}$, $\tau_{c}$ and $\tau_{d}$ are 0.014, 0.02, and 0.06, and 0.1 s respectively (Fig. 2). Suppose the ideal emitted waveform is subject to an elliptic filter with stop and passbands of [3800, 4200] and [3900 , 4100] Hz respectively, a stopband attenuation of 30 dB, and a passband ripple of 1 dB. Also suppose the filter of the receiver’s detector is the same. The peak SNR at the receivers is 50 dB.
With $t=0$, corresponding to the time the switch is thrown, the “ideal” speed of information is the phase speed, $c$ (Fig. 1). For acoustic waves in our scenarios, it equals the phase and group speeds, $c$. The “ideal” time of information is $t_{1}$, and is the time the point of non-analyticity is received if the speed of information is $c$. For application to EM waves (Sec. IV.1), this is the time to beat if the assumption of causality is overturned through observation.
The flat boundary is assumed to have zero pressure fluctuations, and exact solutions are obtained from Eq. 2. The envelopes of the received symbols are derived by taking the absolute value of their Hilbert transforms, both exhibiting their largest value at a time exceeding $t_{1}$ (Fig. 3). As expected, the envelopes of the signals from only the direct paths differ from their temporally-interfering paths and the envelopes all differ after $t_{1}$ (Fig. 4).
Figure 3: Envelopes of received symbol 0 (solid) and 1 (solid). Red line is time the point of non-analyticity arrives at receiver in absence of any finite bandwidth effects along the direct path.
Figure 4: A) Envelope of wave packet for symbol 0 at receiver for direct (black) and temporally interfering paths (green). Red line indicates time when point of non-analyticity arrives if energy propagates at phase and group speed, $c$, in absence of delays from finite-bandwidth effects at source and receiver. B) Same except for symbol 1.
The $c_{3d}$ are derived by cross-correlating their emitted and received timeseries, taking the absolute value of the Hilbert transform, and identifying the lag at the maximum peak, yielding 1694.5 and 2782.5 m/s for symbols 0 and 1 respectively (Fig. 5). Both exceed $c$, thus exhibiting modification of the occurrence of an energy envelope appearing at supersonic speeds in a dispersionless media.
Figure 5: A) Absolute value of Hilbert transform of cross-correlation between emitted and received symbol 0. Ideal time of arrival, $t_{1}$ is vertical dashed line. Red dot at peak. B) Same except for symbol 1 and $t_{1}$ at vertical solid line.
For this example, the BER drops to 0.1 at about 0.005925 s for the temporally-interfering paths (Fig. 6). If the simulation is conducted without a reflecting boundary, only the direct path arrives and the BER drops to 0.1 at 0.0071966 s (Fig. 6). These both occur after $t_{1}$, so the speed of information is less than the phase and group speeds, $c$. Although the peak of the wave packet occurs at supersonic speed, information is transmitted at subsonic speed. Perhaps surprisingly, the speed of information for temporally interfering paths exceeds the speed of information for the direct path only.
Figure 6: BER for symbols 0 & 1 assuming propagation occurs only along direct path (black) or temporally interfering paths (green). tautau axis equals zero at time series sample just before the point of non-analyticity arrives. Red line indicates value of $\tau$ if propagation occurs at phase and group speeds, $c$ and for infinite bandwidths of the switch at the source and the receiver apparatus. BER of 0.1 indicated.
The statistical method for deriving the BER [37] would be unnecessary if the SNR was infinite. Then, one need only check one timeseries point at the receiver, corresponding to when the two infinite SNR timeseries first differ. This is a deterministic decision. One chooses symbol 0 if this first point corresponds to what is received when symbol 0 is transmitted. Otherwise, symbol 1 is declared to have been transmitted. We can mimic almost infinite SNR with a 64 bit precision computer, and implemented this deterministic procedure at a sample frequency of 100 kHz for thousands of random choices for the center frequencies and locations of the source and receiver. All of these cases yielded a speed of information less than the ideal speed, $c$, but not by much. This is not proof the speed of information, $c$, cannot be beaten because the thousands of cases selected is negligible compared to the infinite number of possible symbol waveforms and bandpass filters.
IV Discussion
IV.1 Sound waves and speed of energy and information
Acoustic energy and information can be conveyed at supersonic speeds via shock waves [36]. The absolute value of the Hilbert transform of an acoustic wave packet, derived from the direct+reflected path theory, is a measure of the square root of the energy of sound as a function of time, and thus a measure of energy as a function of time. Therefore, the simulations of the direct+reflected path theory do predict the appearance of supersonic speeds of energy packets but subsonic speed of information in the few simulations conducted so far (Fig. 6). This occurs in a dispersionless media.
Surprisingly, the simulations predict the speed of information for the direct+reflected path effect can exceed the speed for the case with no reflected path at all (Fig. 6). This phenomenon was always observed for the few cases we simulated (not shown). This bears further investigation.
IV.2 Special relativity and the speed of information
The direct+reflected path effect seems to be applicable to waves in general, so we conjecture the appearance of superluminal wave packets may exist for EM waves. An experiment is needed to test this conjecture.
Sommerfeld [60] and Brillouin [6, 7] explain the speed of a sharp discontinuous wavefront of an EM wave moves through media at the speed of light in a vacuum. This sharp wavefront is the so-called point of non-analyticity in more contemporary thinking, and can be replaced by any discontinuity in the emitted waveform. At the time of wavefront passage, they state charges interact with the EM wave, and their re-radiation leads to energy packets with superluminal speeds, but the information flows with the discontinuous wavefront and re-radiation occurs later and therefore cannot increase the speed of information. The experiment conducted by [84] did not find any violation of this interpretation. Their experiment was conducted for superluminal EM wavepackets whose frequencies were near the resonant frequencies of charged particles, though a question arose about the wavepacket’s superluminal speed [50].
We conjecture the direct+reflected path effect could be applicable to EM radiation, and could lead to superluminal wave packets without the presence of charged particles, except within the reflecting boundary. It might be possible to replace the effect of the reflecting boundary with a second source of EM radiation whose radiation mimics the signal along the reflected path, and whose received signal interferes at the receiver with the direct path from the first source. Then there would be no charged particles to re-radiate the EM signals arriving at the receiver, and the physical phenomenon of radiation of excited charges via the Sommerfeld [60], Brillouin [6] mechanism would not be a relevant method to prove information cannot move faster than the speed of light in a vacuum. If it is indeed possible to set up a second source in this way, an outstanding question is whether the speed of information could be made superluminal.
For the direct+reflected path effect, the reflected path’s length always exceeds the direct path’s length. Consequently, with infinite bandwidth filters, the point of non-analyticity has infinite bandwidth and zero temporal resolution, making it impossible for the direct and reflected paths to temporally interfere at the point of non-analyticity when the phase speed is $c$. Thus, the speed of the point of non-analyticity travels at speed $c$, and no faster when only direct and reflected paths exist.
In reality, the point of non-analyticity is made continuous with finite bandwidth filters of bandwidth $\delta f$. These broaden the point to a temporal resolution of $\delta t_{res}=1/\delta f$. Temporal interference occurs if the time between direct and reflected paths is $\delta t_{res}$ or less. Then the broadened point of non-analyticity might be shifted earlier, up to its temporal resolution $\delta t_{res}$, but this earlier shift is subject to the filter delay of $\delta t_{res}$ and it appears no net shift to earlier time occurs. This implies the speed of information does not exceed $c$, and the modified postulate of special relativity is not violated by this physical phenomenon. This is not a proof of non-violation, since the waveform shape following temporal interference differs from the waveform shape passing through the filters, and more distant parts of the waveform may influence the interference pattern near the point of non-analyticity. For the direct+reflected path effect, we conjecture there is a mathematical proof showing the speed of information is less than or equal to $c$ over all classes of functions representing the symbols and bandpass filters because there is no experimental evidence violating the modified postulate of special relativity.
V Conclusion
It is not known if experimental verification of the theory of the direct+reflected path effect would be easier to conduct with acoustic or EM waves, but experimental verification is needed. An optical experiment might be conducted with lasers and acoustic-optic modulators, similar to the apparatus used by Stenner et al. [84], where superluminal speeds are generated by interference derived from temporal interference alone or also with potassium vapors used to increase the group speed due to anomalous dispersion. A beam splitter might send one beam toward the mirror, and the other toward the receiver, where they interfere. If lasers are used to explore the direct+reflected path effect, the amplitudes of the direct and reflected paths would not decay following the spherical spreading of energy, so $a_{i}=1$ in Eq. 2, giving more weight to interference from the reflected path.
We do not understand why the direct+reflected path effect yields a faster speed of information for temporally interfering paths than the speed derived in the absence of the reflected path, where the wave only propagates directly from source to receiver. This may not be true for all types of symbols, and has only been simulated with a few cases. One possibility is interference emphasizes the differences between the symbols just after the point of non-analyticity compared with propagation only along the direct path.
Acknowledgements.
Research was supported by Office of Naval Research grant N00014-23-1-2336.
VI Author Declarations
No conflicts of interest.
The data that support the findings of this study are available within the article.